OpenStudy (anonymous):

what is the slope intercept form of the equation x+3y=-6

OpenStudy (marigirl):

lets change the equation to the gradient and intercept form. as in, make it look like y=mx+c x+3y=-6 3y =-6-x finish it off for me? for y the subject

OpenStudy (marigirl):

slope intercept form is the same as saying gradient and intercept form. sorry forgot to mention that